26. The cost of producing x units of a commodity is given by. TCU) = 100 + 402 – 2x a. Derive the following functions: FC - VNUMBER 26
Answer #1

TC(X) = 100+ nox + 2x2 F = 100 V = Чех нах? AFC - TFC Quantity AVC = TVL - noll 22² a lot 20 ATC = TC = lootuokt 2012 = 100 t

TC(lot)= 2ч, S42 ТcClo-) = loo + keefles ) +2C too) = (eo - це не ооо - гч, тоо TC(Tet) -те ( tse) - гч ѕ42 —гч, оо - ии2 c.

Тc tiн) - TC (x) = мс MC is the additional cost which arises. due to producing an additional mit of output.